OpenStudy (anonymous):

what is equivalent to m/n^7?

jimthompson5910 (jim_thompson5910):

\[\large \frac{m}{n^7}=mn^{-7}\]

OpenStudy (anonymous):

\[m \times 1/ n ^{7 } \rightarrow mn ^{-7}\]